OpenStudy (anonymous):

will give medal!!!!!!!!!!!!!!!!!!!!! Choose the correct answers for amount financed and the finance charge in dollars for the following problem. Curtis Hindle purchased a mountain bike for $650.00 with 75% down. He will make 12 payments of $24.38 each. What is the total amount paid (including down payment) and finance charge in dollars? The total paid (including down payment) is $ . 751.62 780.06 7498.62 The finance charge in dollars is $ . 110.06 130.06 120.06

OpenStudy (anonymous):

@mathmath333 @gorv

OpenStudy (anonymous):

please help @mathmath333

OpenStudy (anonymous):

@KittyT

OpenStudy (anonymous):

@ZnappyDooZ @zpupster @sleepyjess @Synquence

OpenStudy (anonymous):

@RainbowBubbles01

OpenStudy (znappydooz):

Umm... I suck at math :/ I really don't think I'll be able to help u

OpenStudy (anonymous):

ughhh its ok i do to

OpenStudy (znappydooz):

@YanaSidlinskiy is good at it, but she's not online yet

OpenStudy (anonymous):

ok thanks

OpenStudy (zpupster):

$650.00 *.75 =down 12 *24.38= payments payments+ down = total paid total paid -650 = finance charge

OpenStudy (anonymous):

so what the answer @zpupster

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (mathmath333):

zupster has already given u the process, look at it carefully and just simplify the calculations

OpenStudy (anonymous):

did you get it shorty?

OpenStudy (anonymous):

i dont understand it

OpenStudy (anonymous):

:P me niether pfffft

OpenStudy (mathmath333):

\(\large\tt \begin{align*} \color{black}{down=650\times 0.75 \\~\\ total paid=payments+ down \\~\\ total paid =12\times 24.38+650\times 0.75=?\\~\\ finance charge=total paid -650\\~\\ finance charge=(12\times 24.38+650\times 0.75) -650=? }\end{align*}\)